OpenStudy (anonymous):

An adventure game has a number cube with 12 equal-size faces, each labeled with a number from 1 to 12. What is the probability of rolling a 4 or a multiple of 3? Express your answer in simplest form.

OpenStudy (anonymous):

@TheSmartOne @bibby @confluxepic

OpenStudy (bibby):

[1,2,3,4,5,6,7,8,9,10,11,12] which of these is a multiple of 4? a multiple of 3?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

3/12 but in simplest form it would be 1/4?

OpenStudy (bibby):

the or means +

OpenStudy (bibby):

oh I misread it

OpenStudy (bibby):

p(4 or multiple of 3)= p(4)+p(multiple of 3)=1/12+p(multiple of 3)

OpenStudy (bibby):

find the probability of rolling a multiple of 3 (3,6,9,12)